# Interpretable Early Warnings using Machine Learning in an Online Game-experiment

### v1

@@ Abstract
Stemming from physics and later applied to other fields such as ecology, the theory of critical transitions suggests that some regime shifts are preceded by statistical early warning signals. Reddit’s r/place experiment, a large-scale social game, provides a unique opportunity to test these signals consistently across thousands of subsystems undergoing critical transitions. In r/place, millions of users collaboratively created compositions, or pixel-art drawings, in which transitions occur when one composition rapidly replaces another. We develop a machine-learning-based early warning system that combines the predictive power of multiple system-specific time series via gradient-boosted decision trees with memory-retaining features. Our method significantly outperforms standard early warning indicators. Trained on the 2022 r/place data, our algorithm detects half of the transitions occurring within 20 minutes at a false positive rate of just 3.7%. Its performance remains robust when tested on the 2023 r/place event, demonstrating generalizability across different contexts. Using SHapley Additive exPlanations (SHAP) for interpreting the predictions, we investigate the underlying drivers of warnings, which could be relevant to other complex systems, especially online social systems. We reveal an interplay of patterns preceding transitions, such as critical slowing down or speeding up, a lack of innovation or coordination, turbulent histories, and a lack of image complexity. These findings show the potential of machine learning indicators in socio-ecological systems for predicting regime shifts and understanding their dynamics.
This manuscript was compiled on February 14, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
1]0
@@
Abrupt transitions from one state to another are modeled as critical transitions, bifurcations, tipping points, or regime shifts. In social and ecological systems, shifts such as fishery collapses (1, 2), disease outbreaks (3, 4), stock market crashes (5), or the tipping of Earth subsystems (6) directly impact people’s lives. Detecting early signs of critical transitions could enable the relevant actors to prepare for, mitigate, or even avoid such transitions. Yet, finding robust early warning signals is no small task. Empirical data on many transitions in a system is needed to validate models of transitions or learn the patterns preceding them.
@@
Online social systems can offer such detailed observations of many subsystems with intercomparable complex dynamics, and one platform particularly well-suited to studying critical transitions is Reddit’s r/place (7). In this event, which was part game, part social experiment, millions of users created collaborative art by changing the colors of pixels on a shared canvas, inspiring a host of new research to understand the resulting collective behaviors (8, 9, 10, 11, 12, 13, 14, 15). On the canvas, conflict broke out as users fought over the limited space, making r/place analogous to a socio-ecological system. Groups created compositions111In other works discussing r/place, compositions, pixels, and pixel changes are sometimes called artworks, tiles, and updates.—images representing the interests of their community—sometimes by attacking other compositions to rapidly replace their pixels, which we regard as transitions. In this work, we build a warning system for transitions in the social world of r/place, providing a new framework for predicting transitions and understanding their dynamics in large-scale complex systems.
@@
Critical slowing down, the slowing return of the system’s state after perturbations, has been widely investigated as a broadly applicable early warning signal (16, 17, 18, 19, 20). Critical slowing down was shown in ecological experiments more than a decade ago (21, 22, 23), and in theory decades earlier (16). However, its theoretical grounds require the state to closely follow an equilibrium that moves due to a slow parameter change, such as in bifurcations (18), spinodal instabilities (24), continuous or second-order phase transitions, and other reversible, non-catastrophic transitions (25). Critical slowing down may not apply to transitions triggered by a fast parameter change, large perturbations (26), or rate-induced tipping (27); the opposite signal, critical speeding up, even appears in some systems (28, 29, 30). Although other indicators have been proposed (see recent review in Ref. 31), notably based on network (32, 33), dynamical (34), and spatial (35) properties, different warnings have been shown to precede different types of transitions. For some transitions, however, there may be no warnings at all (19, 36), including when long transient dynamics take place before the transition (37). In the absence of a data-validated model, an empirical system cannot be assumed to show any generic warning. Since building empirically sound models is challenging for many complex systems, model-free methods to find early warning indicators are greatly needed.
@@
Early warning indicators leveraging machine learning methods (31, 38) are model-independent and have been shown to outperform generic indicators in some systems (39). Because real-world systems often lack the large amount of data needed for training, some researchers have trained on simulated data (40, 41, 42, 43, 44, 45, 46). However, algorithms using simulated data can only discover predictive features that are embedded in the chosen simulated models, which may not capture all the relevant warning properties of a real-world transition. Models trained on synthetic data and then fine-tuned to empirical data might identify more warning signals (47), but could still overlook features absent in the simulations and too subtle in the sparse tuning data. Another approach is to find systems that provide large amounts of observational data to train on, which has been successful in medicine (48, 49, 50, 51), economics (52, 53) and engineering (54, 55). Nevertheless, techniques that train on large real-world datasets have not yet been applied to predict transitions in socio-ecological systems. Moreover, a standard issue with machine learning predictions is that they rarely provide more understanding of the system to experts or decision makers (56). This “black-box” aspect is not addressed in most studies predicting transitions with machine learning, which focus on the performance of the predictions at the expense of insights into the determinants of the transition.
@@
We use gradient-boosted decision trees to predict the time remaining until impending transitions in Reddit’s r/place, and we analyze the resulting warnings with SHapley Additive exPlanations (SHAP) (57) to understand the mechanisms that drive these transitions. Our warning system predicts transitions well, outperforming each of the generic signals we tested using thresholding and Kendall’s (58). Training on time series of system-specific variables with a 7-hour memory for thousands of compositions, we obtain predictive power up to a few hours before transition. Our predictions are still significant when training on the 2022 event and testing on the 2023 event, which demonstrates that the algorithm exploits genuine and generalizable properties. By examining how different input features contribute to the predictions, we highlight a complex mixture of dynamical properties of compositions near transition, including critical slowing down or speeding up, a lack of innovation or coordination, a turbulent past, and image simplicity. To our knowledge, this analysis provides the first machine learning predictions of critical transitions using detailed observational data in a large-scale complex social system. It reveals drivers of transitions and showcases a framework for building and understanding data-driven warning signals that could be deployed in other socio-ecological scenarios.
@@
To predict transitions, we choose r/place (7) as our study system, as it provides data on complex interactions of millions of participants. Presented as a game without a specific goal, the r/place event allowed registered Reddit users to select any pixel on a canvas, choose a new color from a palette, and then wait for a 5-minute cooldown before changing another pixel (see Fig. 1a). This game-experiment has been hosted by Reddit three times: in 2017, 2022, and 2023. The striking complexity of the resulting collective art is evident in the snapshots of the full 2022 canvas shown in Fig. 1b. Various aspects of the 2017 experiment have been studied (8, 13, 9, 10, 11, 12, 14, 15), while the 2022 and 2023 events have been discussed only once (14).
We use the full data of pixel changes publicly released by Reddit, which includes anonymized user identifiers for each pixel change, for the 2022 and 2023 events. The 2022 event welcomed 10.4 million participants who changed pixel colors 160 million times over 3.5 days, while the 2023 event saw 8.6 million participants making 134 million changes over 5.4 days. We use the 2022 event as our primary system because it was the most popular, with ten times more participants than in 2017, and we use the 2023 event to test the generality of our predictions. The canvas started with 1 million pixels and 16 color choices (8 in 2023) and was later expanded in discrete steps to 4 million pixels (6 million in 2023) with 32 color choices. Detailed rules, statistics, and data provenance on the r/place events are given in section LABEL:secSI:rplacedata of the Supporting Information (SI), along with global canvas distributions in Fig. LABEL:figSI:rplace.
In 2022, users organized to create about 10,900 compositions on the canvas, compared to 6,700 in 2023. These culturally meaningful drawings were often made by communities that previously existed on Reddit or other social media; some also spontaneously formed for the occasion. We identify these compositions using the Atlas dataset (59). This crowdsourced map of the canvas specifies, for each composition, the borders of the image, the birth and death times, the communities involved, and descriptions of the depicted content. We minimally clean the Atlas and separate some compositions so that each is continuous in time and in canvas space (see SI section LABEL:secSI:compodata).
@@
Our definition of a transition in a composition should represent what a user perceives as a sudden change in its image, which translates into a high fraction of differing pixels. Therefore, we define the transition time as the moment this state variable both exceeds 0.35 and is 6 times higher than its average value over the sliding window. Changing these parameters does not significantly affect our predictions, which we show in a sensitivity analysis (SI section LABEL:secSI:sensitivity and Fig. LABEL:figSI:sensitivity). These thresholds ensure that a large portion of the image changes, and that this portion is much larger than usual. We also reject transitions that start from a patchwork of compositions (see examples in Fig. LABEL:figSI:notrans). Because these transitions do not start from a single, coherent composition, this region of the canvas would not be monitored and defended before the new composition exists. Valid transitions include those where the composition returns to its pre-transition state as well as those where a new image takes over (see Fig. 1c for examples of each). More context and justification for how we define transitions is given in SI section LABEL:secSI:transitions. Examples of compositions whose evolving images may resemble transitions but do not pass our criteria are shown in Fig. LABEL:figSI:notrans.
@@
Details, justifications and normalizations for the above variables, as well as additional variables that were implemented but not kept in the training dataset, are discussed in SI section LABEL:secSI:variables. Large correlations might make some variables redundant in the training, which is a criterion to discard a number of variables (see SI section LABEL:secSI:featureselec). The correlation matrix of all the variables used in the training is shown in Fig. LABEL:figSI:correlations. Note that some variables explicitly relate to critical slowing down, such as the variance, autocorrelation, return rate, and duration of last attack; they are used to assess the presence of this generic signal.
@@
We build our warning system by training gradient-boosted decision trees with XGBoost (62) to predict the time-to-transition, . Our training data consists of 1.53 million time instances, each containing feature values that provide predictive information for a given time step and composition. These features consist of the variables listed in the previous section, computed at an instance’s time step or averaged over preceding time periods to emulate a 7-hour memory.
@@
We select time instances to avoid spurious signals. In early versions of this work, we discovered that the transition requirement of a relative threshold on diff pixels reference caused the algorithm to exploit stable pre-transition periods for prediction, rather than using genuine warning signals. Transitions can only be defined on a system in a stable or quasi-stable state, meaning this relative threshold is necessary. To prevent the optimization for stable periods, we require the past sliding-window average of diff pixels reference to be less than for all time instances. This reduces the size of our input dataset by 25%. Other requirements on time instances are explained in SI section LABEL:secSI:filter. Of the 14,160 compositions in 2022, only 6,291 have at least one time instance meeting our requirements, compared to 3,546 compositions out of 6,930 in 2023.
@@
To account for past dynamics, each of 19 time series is recorded with a 7-hour memory for each time instance. This memory consists of 9 or 12 features containing the variable averaged over adjacent time ranges preceding a time instance (see Fig. 3b and SI section LABEL:secSI:memory). The five trivially time-dependent variables are recorded as features without memory. Variables used for training are summarized in Table LABEL:tabSI:variables, and the exact time ranges of the memory are listed in Table LABEL:tabSI:timeranges.
@@
The resulting input data contains 172 features and one target value for 1.53 million time instances in 2022 and 1.34 million in 2023. When using 2022 data only, we attribute 80% of the compositions, with all their instances, to the training dataset and the rest to the test dataset. We also separately train on the full 2022 data to test on the 2023 data. The algorithm outputs a predicted time-to-transition, , for each instance of each test composition (Fig. 3d). The is then calibrated through a monotonic transformation, so that its average is closer to the true time-to-transition, (see SI section LABEL:secSI:calib). Details on the model hyperparameters are given in SI section LABEL:secSI:algo.
@@
We next build binary warnings based on the standard indicators of autocorrelation, return rate, and variance, using two methods. In the first, we use the Kendall’s (58), characterizing the increasing or decreasing trend of a time series in a -minute ( time step) time window. A warning is issued when reaches a threshold, which we sweep over to calculate the ROC and PR curves. In the second method, warnings are issued when the variable itself reaches the threshold.
@@
SHAP values measure, for each instance, the contribution of a certain feature to the target value predicted by the algorithm (57). The distributions of SHAP values over instances for each variable are shown in Fig. LABEL:figSI:shapdistr. When averaged over instances, SHAP absolute values measure feature importance, which we use along with correlations between variables to prune variables and features (SI Section LABEL:secSI:featureselec).
More importantly, SHAP values describe how the algorithm relates a given instance and its feature values to the proximity of transition. Therefore, they can point to characteristics of data that the algorithm associates with incoming transitions. We examine average SHAP values as a function of feature values; these curves for all 172 features are shown in Fig. LABEL:figSI:SHAPall and LABEL:figSI:SHAPnotime. Because some feature distributions are fat-tailed, we use the percentiles of feature values among instances. We interpret the trends of these curves as various warning indicators identified by the algorithm. To ensure that these trends genuinely relate to transition-like dynamics, we only report the trends that are robust against changes in the pre-transition stable period requirement (see SI section LABEL:secSI:shap).

### v2

# pnasresearcharticle Falmagne

@@ Abstract
Stemming from physics and later applied to other fields such as ecology, the theory of critical transitions suggests that some regime shifts are preceded by statistical early warning signals. Reddit’s r/place experiment, a large-scale social game, provides a unique opportunity to test these signals consistently across thousands of subsystems undergoing critical transitions. In r/place, millions of users collaboratively created compositions, or pixel-art drawings, in which transitions occur when one composition rapidly replaces another. We develop a machine-learning-based early warning system that combines the predictive power of multiple system-specific time series via gradient-boosted decision trees with memory-retaining features. Our method significantly outperforms standard early warning indicators. Trained on the 2022 r/place data, our algorithm detects half of the transitions occurring within 20 minutes at a false positive rate of just 3.6%. Its performance remains robust when tested on the 2023 r/place event, demonstrating generalizability across different contexts. Using SHapley Additive exPlanations (SHAP) for interpreting the predictions, we investigate the underlying drivers of warnings, which could be relevant to other complex systems, especially online social systems. We reveal an interplay of patterns preceding transitions, such as critical slowing down or speeding up, a lack of innovation or coordination, turbulent histories, and a lack of image complexity. These findings show the potential of machine learning indicators in socio-ecological systems for predicting regime shifts and understanding their dynamics.
This manuscript was compiled on March 18, 2026 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
@@
Abrupt transitions from one state to another are modeled as critical transitions, bifurcations, tipping points, or regime shifts. In social and ecological systems, shifts such as fishery collapses (jacksonHistoricalOverfishingRecent2001, petrieStructureStabilityExploited2009), disease outbreaks (lagorioQuarantinegeneratedPhaseTransition2011, orozco-fuentesEarlyWarningSignals2019), stock market crashes (vandewalleCrashOctober19871998), or the tipping of Earth subsystems (armstrongmckayExceeding15degCGlobal2022) directly impact people’s lives. Detecting early signs of critical transitions could enable the relevant actors to prepare for, mitigate, or even avoid such transitions. Yet, finding robust early warning signals is no small task. Empirical data on many transitions in a system is needed to validate models of transitions or learn the patterns preceding them.
@@
Online social systems can offer such detailed observations of many subsystems with intercomparable complex dynamics, and one platform particularly well-suited to studying critical transitions is Reddit’s r/place (Place2024). In this event, which was part game, part social experiment, millions of users created collaborative art by changing the colors of pixels on a shared canvas, inspiring a host of new research to understand the resulting collective behaviors (armstrongCoordinationPeerProduction2018, rappazLatentStructureCollaboration2018, vachherUnderstandingCommunityLevelConflicts2020, litherlandInstructionVsEmergence2021, israeliFlyingColorsPredicting2023, mullerCompressionCulturalEvolution2018, wuLargescaleCollectiveDynamics2024, botelhoArtExpandedField2024). On the canvas, conflict broke out as users fought over the limited space, making r/place analogous to a socio-ecological system. Groups created compositions111In other works discussing r/place, compositions, pixels, and pixel changes are sometimes called artworks, tiles, and updates.—images representing the interests of their community—sometimes by attacking other compositions to rapidly replace their pixels, which we regard as transitions. In this work, we build a warning system for transitions in the social world of r/place, providing a new framework for predicting transitions and understanding their dynamics in large-scale complex systems.
@@
Critical slowing down, the slowing return of the system’s state after perturbations, has been widely investigated as a broadly applicable early warning signal (wisselUniversalLawCharacteristic1984, vannesSlowRecoveryPerturbations2007, schefferEarlywarningSignalsCritical2009, boettigerEarlyWarningSignals2013, boettigerPatternsPredictions2013). Critical slowing down was shown in ecological experiments more than a decade ago (carpenterEarlyWarningsRegime2011, daiGenericIndicatorsLoss2012, dakosMethodsDetectingEarly2012), and in theory decades earlier (wisselUniversalLawCharacteristic1984). However, its theoretical grounds require the state to closely follow an equilibrium that moves due to a slow parameter change, such as in bifurcations (schefferEarlywarningSignalsCritical2009), spinodal instabilities (levinPhaseTransitionsTheory2023), continuous or second-order phase transitions, and other reversible, non-catastrophic transitions (kefiEarlyWarningSignals2013). Critical slowing down may not apply to transitions triggered by a fast parameter change, large perturbations (boettigerNoEarlyWarning2013), or rate-induced tipping (ritchieEarlywarningIndicatorsRateinduced2016); the opposite signal, critical speeding up, even appears in some systems (titusCriticalSpeedingEarly2020, pomeauCriticalSpeedvsCritical2011, gietkaSqueezingCriticalSpeeding2022). Although other indicators have been proposed (see recent review in Ref. georgeEarlyWarningSignals2023), notably based on network (wunderlingHowMotifsCondition2020, suweisEarlyWarningSigns2014), dynamical (xuNonequilibriumEarlywarningSignals2023), and spatial (chenEigenvaluesCovarianceMatrix2019) properties, different warnings have been shown to precede different types of transitions. For some transitions, however, there may be no warnings at all (boettigerEarlyWarningSignals2013, schefferAnticipatingCriticalTransitions2012), including when long transient dynamics take place before the transition (hastingsTransientPhenomenaEcology2018). In the absence of a data-validated model, an empirical system cannot be assumed to show any generic warning. Since building empirically sound models is challenging for many complex systems, model-free methods to find early warning indicators are greatly needed.
@@
Early warning indicators leveraging machine learning methods (georgeEarlyWarningSignals2023, obrienEWSmethodsPackageForecast2023) are model-independent and have been shown to outperform generic indicators in some systems (choiEarlyWarningCritical2022). Because real-world systems often lack the large amount of data needed for training, some researchers have trained on simulated data (qiUsingMachineLearning2020, brettDynamicalFootprintsEnable2020, grassiaMachineLearningDismantling2021, buryDeepLearningEarly2021, debMachineLearningMethods2022, dylewskyUniversalEarlyWarning2023, miryDeepLearningDisease2025). However, algorithms using simulated data can only discover predictive features that are embedded in the chosen simulated models, which may not capture all the relevant warning properties of a real-world transition. Models trained on synthetic data and then fine-tuned to empirical data might identify more warning signals (liuEarlyPredictorOnset2024), but could still overlook features absent in the simulations and too subtle in the sparse tuning data. Another approach is to find systems that provide large amounts of observational data to train on, which has been successful in medicine (tapakComparativeEvaluationTime2019, gaoMachineLearningBased2020, hylandEarlyPredictionCirculatory2020, kobylarzribeiroMachineLearningEarly2020), economics (samitasMachineLearningEarly2020, barthelemyEarlyWarningSystem2024) and engineering (maDatadrivenPowerSystem2018, lassetterUsingCriticalSlowing2021). Nevertheless, techniques that train on large real-world datasets have not yet been applied to predict transitions in socio-ecological systems. Moreover, a standard issue with machine learning predictions is that they rarely provide more understanding of the system to experts or decision makers (rudinStopExplainingBlack2019). This “black-box” aspect is not addressed in most studies predicting transitions with machine learning, which focus on the performance of the predictions at the expense of insights into the determinants of the transition.
@@
We use gradient-boosted decision trees to predict the time remaining until impending transitions in Reddit’s r/place; this time is set to 12 h when there is no incoming transition or it is far in the future. Critically, we analyze the resulting warnings with SHapley Additive exPlanations (SHAP) (lundbergUnifiedApproachInterpreting2017) to understand the mechanisms that drive these transitions. Our warning system predicts transitions well, outperforming each of the generic signals we tested using thresholding and Kendall’s (chenPracticalGuideUsing2022). Training on time series of system-specific variables with a 7-hour memory for thousands of compositions, we obtain predictive power up to a few hours before transition. Our predictions are still significant when training on the 2022 event and testing on the 2023 event, which demonstrates that the algorithm exploits genuine and generalizable properties. By examining how different input features contribute to the predictions, we highlight a complex mixture of dynamical properties of compositions near transition, including critical slowing down or speeding up, a lack of innovation or coordination, a turbulent past, and image simplicity. To our knowledge, this analysis provides the first machine learning predictions of critical transitions using detailed observational data in a large-scale complex social system. It reveals drivers of transitions and showcases a framework for building and understanding data-driven warning signals that could be deployed in other socio-ecological scenarios.
@@
To predict transitions, we choose r/place (Place2024) as our study system, as it provides data on complex interactions of millions of participants. Presented as a game without a specific goal, r/place allowed registered Reddit users to select any pixel on a canvas, choose a new color from a palette, and then wait for a 5-minute cooldown before changing another pixel (see Fig. 1a). We use the term event to refer to each r/place game-experiment that Reddit hosted, which happened in 2017, 2022 and 2023. Each event lasted several days and followed the same basic rules. The striking complexity of the resulting collective art is evident in the snapshots of the full 2022 canvas shown in Fig. 1b. Various aspects of the 2017 experiment have been studied (armstrongCoordinationPeerProduction2018, mullerCompressionCulturalEvolution2018, rappazLatentStructureCollaboration2018, vachherUnderstandingCommunityLevelConflicts2020, litherlandInstructionVsEmergence2021, israeliFlyingColorsPredicting2023, wuLargescaleCollectiveDynamics2024, botelhoArtExpandedField2024), while the 2022 and 2023 events have been discussed only once (wuLargescaleCollectiveDynamics2024).
We use the full data of pixel changes publicly released by Reddit, which includes anonymized user identifiers for each pixel change, for the 2022 and 2023 events. The 2022 event welcomed 10.4 million participants who changed pixel colors 160 million times over 3.5 days, while the 2023 event saw 8.6 million participants making 134 million changes over 5.4 days. We use the 2022 event as our primary system because it was the most popular, with ten times more participants than in 2017, and we use the 2023 event to test the generality of our predictions. The canvas started with 1 million pixels and 16 color choices (8 in 2023) and was later expanded in discrete steps to 4 million pixels (6 million in 2023) with 32 color choices—see SI Table LABEL:tabSI:extensions. Detailed rules, statistics, and data provenance on the r/place events are given in Section LABEL:secSI:rplacedata of the Supporting Information (SI), along with global canvas distributions in Fig. LABEL:figSI:rplace.
In 2022, users organized to create about 10,900 compositions on the canvas, compared to 6,700 in 2023. These culturally meaningful drawings were often made by communities that previously existed on Reddit or other social media; some also spontaneously formed for the occasion. We identify these compositions using the Atlas dataset (haagmansPlaceAtlasInitiative2024). This crowdsourced map of the canvas specifies, for each composition, the borders of the image, the birth and death times, the communities involved, and descriptions of the depicted content. We minimally clean the Atlas and separate some compositions so that each is continuous in time and in canvas space (see SI Section LABEL:secSI:compodata).
@@
Our definition of a transition in a composition represents what a user perceives as a sudden change in its image, which translates into a high fraction of differing pixels. Therefore, we define the transition time as the moment this state variable both exceeds 0.35 and is 6 times higher than its average value over the sliding window. Changing these parameters, including the sliding window length, does not substantially affect our performance, which we show in a sensitivity analysis (SI Section LABEL:secSI:sensitivity and Fig. LABEL:figSI:sensitivitya). These thresholds ensure that a large portion of the image changes, and that this portion is much larger than usual. The relative threshold of 6 selects for sudden transitions, which are harder to predict and of more interest to the warning signals literature, while requiring a stable system to transition from. We ran an alternative training with only a relative threshold of 2: this effectively includes smoother transitions, and performance is increased (see Fig. LABEL:figSI:sensitivityc), but these transitions are harder to interpret and less relevant to real-world warnings. We also reject transitions that start from a patchwork of compositions (see examples in Fig. LABEL:figSI:notrans). Because these transitions do not start from a single, coherent composition, this region of the canvas would not be monitored and defended before the new composition exists. Valid transitions include those where the composition returns to its pre-transition state as well as those where a new image takes over (see Fig. 1c for examples of each). More context and justification for how we define transitions is given in SI Section LABEL:secSI:transitions. Examples of compositions whose evolving images may resemble transitions but do not pass our criteria are shown in Fig. LABEL:figSI:notrans.
@@
(Fig. 2r). Change relative to the sliding window of an entropy proxy, the ratio of the compressed file size of the image to the image area, as introduced in Ref. martinianiQuantifyingHiddenOrder2019.
@@
(Fig. 2s). Change relative to the sliding window of the fractal dimension, calculated using the reticular box counting method separated by color as described in Ref. bisoiCalculationFractalDimension2001.
@@
Details, justifications and normalizations for the above variables are in SI Section LABEL:secSI:variables. Note that some variables explicitly relate to critical slowing down, such as the variance, autocorrelation, return rate, and duration of last attack; they are used to assess the presence of this generic signal.
Large correlations make some variables redundant in the training, which is a criterion to discard a number of variables (see SI Section LABEL:secSI:featureselec). The correlation matrix of all the variables used in the training is shown in Fig. LABEL:figSI:correlations. The variables that were implemented but rejected from the final training dataset are also described in SI Section LABEL:secSI:variables. A subset of these variables is related to additional spatial properties of the image dynamics: levenshtein complexity, multiscale complexity, and those described in SI Section LABEL:secSI:spatialvars (SI Table LABEL:tabSI:additionalvariables after pruning); these variables are included in an extended set of variables on which we train the algorithm. While these variables individually have predictive power (see Fig. LABEL:figSI:shapdistrb and c), including them does not improve on the total performance of the nominal training, even after pruning the added variables as detailed in SI Section LABEL:secSI:featureselec (see Fig. LABEL:figSI:sensitivityb). Since they also add complexity—and thus decrease interpretability—these variables are excluded from the training.
@@
We build our warning system by training gradient-boosted decision trees with XGBoost (chenXGBoostScalableTree2016) to predict the time-to-transition, . Our training data consists of 1.53 million time instances, each containing feature values that provide predictive information for a given time step and composition. These features consist of the variables listed in the previous section, computed at an instance’s time step or averaged over preceding time periods to emulate a 7-hour memory.
@@
We select time instances to avoid spurious signals. In early versions of this work, we discovered that the transition requirement of a relative threshold on diff pixels reference caused the algorithm to exploit stable pre-transition periods for prediction, rather than using genuine warning signals. Transitions can only be defined on a system in a stable or quasi-stable state, meaning this relative threshold is necessary. To prevent the optimization for stable periods, we require the past sliding-window average of diff pixels reference to be less than for all time instances. The effect of removing this filter on the performance is shown in the “no stable times” curve of Fig. LABEL:figSI:sensitivityc. This reduces the size of our input dataset by a third. Other requirements on time instances are explained in SI Section LABEL:secSI:filter. Of the 14,160 compositions in 2022, only 6,289 have at least one time instance meeting our requirements, compared to 3,546 compositions out of 6,930 in 2023.
@@
To account for past dynamics, each of 19 time series is recorded with a 7-hour memory for each time instance. This memory consists of 9 or 12 features (see SI Section LABEL:secSI:memory for explanation of the number of time features) containing the variable averaged over adjacent time ranges preceding a time instance (see Fig. 3b and SI Section LABEL:secSI:memory). The five trivially time-dependent variables are recorded as features without memory. Variables used for training are summarized in SI Table LABEL:tabSI:variables, and the exact time ranges of the memory are listed in SI Table LABEL:tabSI:timeranges.
@@
The resulting input data contains 175 features and one target value for 1.53 million time instances in 2022 and 1.34 million in 2023. When using 2022 data only, we attribute 80% of the compositions, with all their instances, to the training dataset and the rest to the test dataset. We also separately train on the full 2022 data to test on the 2023 data. The algorithm outputs a predicted time-to-transition, , for each instance of each test composition (Fig. 3d). The is then calibrated through a monotonic transformation, so that its average is closer to the true time-to-transition, (see SI Section LABEL:secSI:calib). Details on the model hyperparameters are given in SI Section LABEL:secSI:algo.
@@
We next build binary warnings based on the standard indicators of autocorrelation, return rate, and variance, using two methods. In the first, we use the Kendall’s (chenPracticalGuideUsing2022), characterizing the increasing or decreasing trend of a time series in a -minute ( time step) time window. A warning is issued when reaches a threshold, which we sweep over to calculate the ROC and PR curves. In the second method, warnings are issued when the variable itself reaches the threshold.
@@
SHAP values measure, for each instance, the contribution of a certain feature to the target value predicted by the algorithm (lundbergUnifiedApproachInterpreting2017). The distributions of SHAP values over instances for each variable are shown in Fig. LABEL:figSI:shapdistr (panel (a) for the nominal variable selection). When averaged over instances, SHAP absolute values measure feature importance. We use this along with correlations between the features values and correlations between their SHAP values to prune variables and features (SI Section LABEL:secSI:featureselec).
Importantly, SHAP values describe how the algorithm relates a given instance and its feature values to the proximity of transition. Therefore, they can point to characteristics of data that the algorithm associates with incoming transitions. We examine average SHAP values as a function of feature values; these curves for all 175 features are shown in Fig. LABEL:figSI:SHAPall and LABEL:figSI:SHAPnotime. Because some feature distributions are fat-tailed, we use the percentiles of feature values among instances. We interpret the trends of these curves as various warning indicators identified by the algorithm. To ensure that these trends genuinely relate to transition-like dynamics, we only report the trends that are robust against changes in the pre-transition stable period requirement (see SI Section LABEL:secSI:shap).